(ii) 2/√x + 3/√y = 2

4/√x + 9/√y = -1

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Solution: Substituting 1/√x = m and 1/√y = n in the given equations, we get 2m + 3n = 2 ………………………..(i) 4m – 9n = -1 ………………………(ii) Multiplying equation (i) by 3, we get 6m + 9n = 6 ………………….…..(iii) Adding equation (ii) and (iii), we get 10m = 5 m = 1/2…………………………….…(iv) Now by putting the value of ‘m’ in equation (i), we get 2×1/2 + 3n = 2 3n = 1 n = 1/3 m =1/√x ½ = 1/√x x = 4 n = 1/√y 1/3 = 1/√y y = 9 Hence, x = 4 and y = 9